OpenStudy (anonymous):

What is the value of time, t, for the given values of distance, d, and rate, r? d = 425 m, r = 50 m/s A. 8.5 m B. 8.5 s C. 21,250 m D. 21,250 s

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Would it make sense to say, that \(\large\color{slate}{ d=t\times r }\) ? Saying that for example, you walked 4 hours, at a rate of 3 mile per hour, then your distance would be: \(\large\color{slate}{ d=t\times r }\) \(\large\color{slate}{ d=5(hours)\times 3(miles/hours) }\) (the hours cancel, and your unit is:) \(\large\color{slate}{ d=15(miles) }\)

OpenStudy (solomonzelman):

I mean you walked \(\large\color{red}{ 5}\) hours, at a rate of 3 miles per hour....

OpenStudy (solomonzelman):

Is this example making sense?

OpenStudy (anonymous):

no not at all im so confused by this all

OpenStudy (solomonzelman):

OK, lets say `Sam ran 2 hour, with a speed of 7 miles per hour. ` `Would you be able to tell me the distance Sam ran? `

OpenStudy (anonymous):

yes 7 + 7 = 14 so sam ran 14 miles right?

OpenStudy (solomonzelman):

but why not multiply 7 times 2?

OpenStudy (solomonzelman):

would you be adding, if Sam ran 100 miles?

OpenStudy (anonymous):

no probly not

OpenStudy (solomonzelman):

yes, so if he ran for 100 hours, at a speed of 7 miles per hour, what would you do to calculate the distance ran?

OpenStudy (anonymous):

7 x 100?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So now if you know that Sam ran 800 miles, and he was running at a speed of 4 miles per hour, how would you find the number of hours he ran for?

OpenStudy (anonymous):

4 x 800?

OpenStudy (solomonzelman):

no, you already given the distance, it is 800 miles, and he ran with a speed of 4 miles per hour.

OpenStudy (solomonzelman):

Wouldn't you divide the distance (i.e. 800 miles) by the rate (i.e. 4 miles per hour) ?

OpenStudy (anonymous):

ah yes

OpenStudy (solomonzelman):

so, you have the same situtation here

OpenStudy (solomonzelman):

Sam ran 425 meters, at a rate of 50 meters per second.

OpenStudy (solomonzelman):

How much did time did he run for?

OpenStudy (anonymous):

425 divid by 50?

OpenStudy (solomonzelman):

yes, there you go!

OpenStudy (solomonzelman):

your answer is?

OpenStudy (anonymous):

8.5 or B

OpenStudy (solomonzelman):

What did we find just now to be 8.5, the distance traveled or the time?

OpenStudy (solomonzelman):

what have we *found*

OpenStudy (anonymous):

the time right?

OpenStudy (solomonzelman):

Yes,

OpenStudy (solomonzelman):

So your answer is?

OpenStudy (solomonzelman):

A or B ?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

wait A

OpenStudy (anonymous):

its A...........right?

OpenStudy (solomonzelman):

B, it is seconds, --- time

OpenStudy (anonymous):

oh i see

OpenStudy (solomonzelman):

it is not in meters, b/c that would be distance